OpenStudy (anonymous):

Suppose you need to install two heaters, which will be powered by conventional residential power, i.e. V=120. V, and are to be controlled by a single on-off switch. Each heater works by dissipating power in a resistor. The first heater, with resistance R1, is to produce P1=560. W. The second, with resistance R2, is to produce P2=800. W. a) If the two heaters are to be wired in series, what should R1 and R2 be? b) If the two heaters are to be wired in parallel, what should R1 and R2 be?

OpenStudy (anonymous):

I think i did this question lol, where do you go?

OpenStudy (anonymous):

Ohh :/ no lol, I'm in H.S D;

OpenStudy (anonymous):

In parallel, the voltage across each resistor is the same and the power is V^2/R1 and V^2/R2 In series, the current is the same for both, and the power is I^2 R1 and I^2 R2, with the current I being I = V(R1+R2).

OpenStudy (anonymous):

No, I meant in series the current is I = V/(R1+R2).

OpenStudy (radar):

For the series arrangement first calculate the current required for the power required. Power required: 560 watt plus 800 watts = total of 1360 watts. Voltage to be used IAW problem specification is 120 VAC. Resultant current needed is P = EI, 1360=120I, I then is 11 1/3 Amps. Knowing the current, you now can calculate R1 and R2 has you have the two powers required (see problem specs) use P=I^2R for those two powers getting R1 and R2.

OpenStudy (radar):

For the parallel arrangement use the two powers required and the AC voltage 120 VAC P = (E^2)/R solving for R1 and R2.

OpenStudy (anonymous):

Thanks for the help!!